2. Find the relation between x and y if the points (x, y), (1, 2) and (7, 0) are collinear.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
If given points are collinear then area of triangle formed by them must be zero. Let (x, y), (1, 2) and (7, 0) are vertices of a triangle, Area of a triangle = 1/2 × [x1(y2 – y3) + x2(y3 – y1) + x3(y1 – y2)] = 0 [x(2 – 0) + 1 (0 – y) + 7( y – 2)] = 0 2x – y + 7y – 14 = 0 2x + 6y – 14 = 0 x + 3y – 7 = 0. Which is the required result.